table 2.3 Corporate Tax Rates as of 2015 Pay this Amount on Base Income Plus this Percentage on Anything Over the Base Taxable Income $0-$50,000 $50,001-$75,000 $75,001-$100,000 $100,001-$335,000 $335,001-$10,000,000 $10,000,001-$15,000,000 $15,000,001-$18,333,333 Over $18,333,333 0 7,500 13,750 22,250 113,900 3,400,000 5,150,000 6,416,667 15% 25 34 39 34 35 38 35Hunt Taxidermy, Inc. is concerned about the taxes paid by the company in 2018. In addition to $4.4 million of taxable income, the firm received $309,000 of interest on state-issued bonds and $240,000 of dividends on common stock it owns in Oakdale Fashions, Inc (Use Table 2.3.) Calculate Hunt Taxidermys tax liability. (Enter your answer in dollars not in millions. Round your answer to the nearest dollar amount.) Tax liability Calculate Hunt Taxidermys average tax rate. (Round your answer to 2 decimal places.) Average tax rate Calculate Hunt Taxidermys marginal tax rate. Marginal tax rate
Answer #1

Answer:

Given

Taxable income=$4.4 million

Interest Income =$309,000=$0.309 millions

Dividend Income=$240000=$0.24 millions

So Total Taxable income=Taxable income+Interest Income+Dividend Income=4.4+0.309+0.24=$ 4.949

A) Since taxable income is between $335001-10,000,000

So Tax Laibility =0.113900+34%*(4.949-0.335000)=$1.68266 millions

B) Average tax rate= Tax liability / Total Taxable income=1.68266/4.949=0.34=34%

C) Since taxable income is between $335001-10,000,000 so marginal tax rate =34%